# A user-friendly SPARQL query editor powered by lightweight metadata

### v1

@@ 2.2 A context aware autocomplete
Our editor retrieves the VoID metadata, if available, by querying directly the endpoint itself. The main information of interest processed by the editor from the VoID descriptions is represented by the classes instantiated in the endpoint and properties that relate them. This ensures that only properties that are stated and applicable according to the schema and the known type of a variable are suggested in the autocomplete drop-down menu (see Figure 1). If no VoID metadata are found, the editor will try to retrieve all classes and properties. Finally, our autocomplete approach is lighter, faster, and complementary when compared to the state-of-the-art approaches such the Qlever autocomplete. In addition, it works within SPARQL 1.1 federated SERVICE calls.
@@ 3 Demonstration
The proposed SPARQL query editor is publicly available and deployed at https://sib-swiss.github.io/sparql-editor. Users can choose either a predefined endpoint shown in the SPARQL endpoint field at the top, or input their own endpoint of interest in this field. Note that the new features described in Section 2 will only be available if the corresponding metadata are accessible through the assigned SPARQL endpoint. At the time of writing, 8 out of 10 public endpoints listed in our official query editor release included all the metadata needed for our tool to be fully operational. This also demonstrates that our approach is triple store agnostic, and readily deployable over diverse RDF data sources.
@@ 4 Conclusion
We have presented an easy-to-deploy, open source query editor over any SPARQL endpoint, which includes several novel features poweredby lightweight metadata. The proposed features can greatly help SPARQL-savvy users to write queries over any SPARQL endpoint of interest. We also successfully applied our query editor to several public SPARQL endpoints that range from small to very large real-word RDF datasets. Finally, we welcome contributions from the community to develop this further in our GitHub repository.
@@ 4.0.1 Acknowledgements
CHIST-ERA-22-ORD/SNF grant 20CH21_217482; the swissuniversities CHORD in Open Science I, grant “Swiss DBGI-KM”.

### v2

@@ 2.2 A context aware autocomplete
Our editor retrieves the VoID metadata, if available, by querying directly the endpoint itself. The main information of interest processed by the editor from the VoID descriptions is represented by the classes instantiated in the endpoint and properties that relate them. This ensures that only properties that are stated and applicable according to the schema and the known type of a variable are suggested in the autocomplete drop-down menu (see Figure 1). If no VoID metadata are found, the editor will try to retrieve all classes and properties. Finally, our autocomplete approach is lighter, faster, and complementary when compared to the state-of-the-art approaches such as the Qlever autocomplete. In addition, it works within SPARQL 1.1 federated SERVICE calls.
@@ 3 Demonstration
The proposed SPARQL query editor is publicly available and deployed at https://sib-swiss.github.io/sparql-editor. Users can choose either a predefined endpoint shown in the SPARQL endpoint field at the top, or input their own endpoint of interest in this field. Note that the new features described in Section 2 will only be available if the corresponding metadata are accessible through the assigned SPARQL endpoint. At the time of writing, 8 out of 10 public endpoints listed in our official query editor release included all the metadata needed for our tool to be fully operational. This also demonstrates that our approach is triple store agnostic, and readily deployable over diverse RDF data sources. Finally, non-savvy SPARQL users can use our ExpasyGPT tool at https://www.expasy.org/chat that generates SPARQL queries from plain-English text and run these queries with our SPARQL editor.
4 Limitations
  • Editor features limitations: YASGUI relies on CodeMirror 5101010https://codemirror.net/5/ which is outdated and does not have all the latest features of CodeMirror 6. We investigated the later one, but we realised that upgrading it would require to rewrite several core functionalities, such as autocomplete. Alternatively, we also consider replacing the editor with a VisualStudio based editor and a language server protocol (LSP) for SPARQL, for example, the one developed by the Qlever team [^11] https://github.com/IoannisNezis/Qlue-ls.
  • Property autocomplete: For property autocomplete to fully work, users need to explicitly define the subject type. This could be improved by reusing methods from the efficient and effective SPARQL autocompletion paper[1].
  • Target users: This tool is intended for users that knows the SPARQL query language. Nevertheless, writing a query often requires exploratory steps to understand the data schema, that is how the data are structured. Our editor aims to reduce this friction and assist users in dealing with complex or unfamiliar endpoints.
  • Visual query builder: Supporting less tech-savvy users through visual query building is currently out of scope; we recognize its value and believe our work on exposing and standardizing endpoint metadata could serve as a foundation for such tools. We welcome future contributions or collaborations that aim to expand the tool in this direction.
@@ 5 Conclusion
We have presented an easy-to-deploy, open source query editor over any SPARQL endpoint, which includes several novel features powered by lightweight metadata. The proposed features can greatly help SPARQL-savvy users to write queries over any SPARQL endpoint of interest. We also successfully applied our query editor to several public SPARQL endpoints that range from small to very large real-word RDF datasets. As a future work, we aim at improving the autocomplete feature to suggest next triple patterns by also considering predefined functions and rules. Currently, the proposed features will not work properly, if the needed metadata are not provided. To mitigate the later issue, we will work on an approach that probes endpoints and RDF data dumps, if available. Finally, we welcome contributions from the community to develop this further in our GitHub repository.
@@ 5.0.1 Acknowledgements
This work was funded by the Swiss National Science Foundation (SNSF) under the CHIST-ERA-22-ORD/SNSF TRIPLE project, grant: 20CH21_217482; the SNSF MetaboLinkAI project, grant: 10.002.786; and the swissuniversities CHORD in Open Science I, grant: “Swiss DBGI-KM”.